OpenStudy (anonymous):

(a+5)(a-3) find the product

OpenStudy (texaschic101):

FOIL method (first, outside, inside, last) (a + 5)(a - 3) F -- multiply first terms in each set a * a = a^2 O -- multiply outside terms in each set a * -3 = -3a I --- multiply inside terms in each set 5 * a = 5a L --- multiply last terms in each set 5 * -3 = -15 now we have : a^2 - 3a + 5a - 15 -- combine like terms a^2 + 2a - 15 any questions ?

OpenStudy (texaschic101):

yes....

OpenStudy (texaschic101):

I can try